The minute hand of a clock is 7cm long. to the nearest tenth of a centimeter, how far does the tip of the minute hand move in 25 minutes ?
Found 2 solutions by stanbon, nerdybill:
Answer by stanbon(75887)

The minute hand of a clock is 7cm long. to the nearest tenth of a centimeter, how far does the tip of the minute hand move in 25 minutes
------------
Circumference = (pi)d = (pi)14 cm
------------------------
Arc of the minute hand = (25/60)(14(pi)) cm
Arc length = 5.833333(pi) = 18.33 cm
========================================
Cheers,
Stan H.

Answer by nerdybill(7384)

In an hour, the minute hand moves 360 degrees -- full circle around the clock.
The total distance in an hour is equal to the circumference:
(pi)(2r)
= (3.14)(2)7
= (3.14)(14)
= 43.96 cm (distance if it traveled 60 minutes)
.
If it only moved 25 minutes:
25/60 * 43.96
= 18.317 cm